OpenStudy (anonymous):

An 80.0-gram sample of a gas was heated from 25 °C to 225 °C. During this process, 346 J of work was done by the system and its internal energy increased by 7905 J. What is the specific heat of the gas?

OpenStudy (aaronq):

internal energy is given by: U = q + w the heat can be found with: \(q=m*C_p*\Delta T\) substituting (2) into (1): \(U = (m*C_p*\Delta T) + w\)

OpenStudy (aaronq):

solve for specific heat capacity \((C_p)\).

OpenStudy (anonymous):

I'm not sure of what to put where..I know w is -346 and m 80

OpenStudy (aaronq):

"internal energy increased by 7905 J" \(\Delta U=7905 J=m*C_p*(T_f-T_i)+w\) because \(\Delta T=T_f-T_i\)

OpenStudy (anonymous):

I get .505 but I think that is wrong

OpenStudy (anonymous):

so I got 7905J=80g*Cp*200+-346 but .505 is wrong

OpenStudy (aaronq):

i got Cp=0.5156875

OpenStudy (anonymous):

that is the correct answer but how did you get that...also thank you

OpenStudy (aaronq):

No problem, i used that same equation, you probably messed up in the arithmetic because you were pretty close.